News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 20 June 2003 7.30am EDT

During the last week Chandra completed the observing schedule as planned.

Following the warm-restart of IRU-1 gyro-1 on 13 June, its current has returned from the restart value of ~100 mA to approximately the same level as prior to the warm-restart (~140 mA). This behavior is consistent with a trend towards eventual out of spec (anomalous) performance and the team is currently assessing a swap to IRU-2 in order to retain maximum operational lifetime on gyro-1. IRU-2 has continued to operate in parallel with IRU-1 and the initial calibration was completed last week without impact to operations.

A Chandra press release was issued on June 19 describing results from the Great Observatories Origins Deep Survey (GOODS) that combines observations from both Chandra and Hubble. For details see:

The observing schedule for next week is shown below and includes a follow-up observation of the Target of Opportunity H1743-322.

Radiation Belts                     Jun 22
H1743-322              ACIS-S/HETG  Jun 23
Mrk 1044                            Jun 24
SDSS J014812.23+0001   ACIS-S       Jun 25
Radiation Belts
AX J1845-0258          ACIS-S       Jun 26
AX J1845-0258          ACIS-S
G11.2-0.3              ACIS-S       Jun 27
NGC 6440               ACIS-S
NGC 5728               ACIS-S
DLS J0522-4820         ACIS-I
Radiation Belts                     Jun 28
AX J1845-0258          ACIS-S
Markarian 290          ACIS-S/HETG


All spacecraft subsystems, with the exception of the MUPS thruster A-1 under investigation, continued to operate nominally.

Return to Status Reports